# Understanding World or Predicting Future? A Comprehensive Survey of World Models

### v2

@@ 6.5. Ethical and Safety Concerns
6.6. Benchmark
Benchmarking world models is both necessary and challenging. Because the community pursues divergent goals—learning internal representations vs. predicting future worlds—with heterogeneous technical approaches (e.g., LLM agents, video diffusion) and wide-ranging application domains (autonomous driving, robotics, social simulation), there is no single canonical task or metric. Nevertheless, several recent efforts illustrate how carefully designed testbeds can expose the specific gaps that prevent current models from becoming reliable world simulators.
Video-centric world simulation. Qin et al. propose WorldSimBench(qin2024worldsimbench, ), a dual evaluation suite combining human-preference video judgments and action-level consistency across three embodied settings (open-ended sandboxes, autonomous driving, and robot manipulation). Complementarily, Duan et al. present WorldScore(duan2025worldscore, ), which decomposes “world generation” into controllability, visual quality, and dynamics across 3,000 camera-specified scenes, enabling head-to-head comparisons of 3D, 4D, and video generators.
Physical and spatial reasoning. PhysBench (chow2025physbench, ) (10k video–image–text triplets) and UrbanVideo-Bench (zhao2025urbanvideo, ) (5.2k drone clips) demonstrate that current VLMs lack basic physics understanding and urban navigation skills. Xu et al. translate psychometric testing into AI evaluation by defining five Basic Spatial Abilities and diagnosing geometry and rotation weaknesses across 13 VLMs (xu2025defining, ).
Embodied decision making. The Embodied Agent Interface (EAI) (li2024embodied, ) standardizes four LLM-based modules—goal parsing, sub-goal generation, action sequencing, and transition modeling—and reports fine-grained errors rather than a single success rate.
@@ 7. Conclusion
Understanding the world and predicting the future have been long-standing objectives for scientists developing artificial generative intelligence, underscoring the significance of constructing world models across various domains. This paper presents the first comprehensive survey of world models that systematically explores their two primary functionalities: implicit representations and future predictions of the external world. We provide an extensive summary of existing research on these core functions, with particular emphasis on world models in decision-making, world knowledge learned by models, world models as video generation, and world models as embodied environments. Additionally, we review progress in key applications of world models, including autonomous driving, robotics, and social simulation. Finally, recognizing the unresolved challenges in this rapidly evolving field, we highlight open problems and propose promising research directions with the hope of stimulating further investigation in this burgeoning area.

### v4

@@ 6.2. Enriching the Social Dimension
6.3. Benchmarks
Table 8. Representative benchmarks for evaluating world models.
Category Benchmark Scope & highlights
Video-centric world simulation WorldSimBench (qin2024worldsimbench, ) Sandbox/driving/manipulation; links human preference to action-level consistency
WorldScore (duan2025worldscore, ) 3,000 camera-specified scenes; decomposes controllability/quality/dynamics; compares 3D/4D/video generators
VBench (huang2024vbench, ) General T2V/V2V; automatic axes: temporal consistency, stability, prompt adherence
VBench-2.0 (zheng2025vbench, ) General T2V/V2V; intrinsic faithfulness: physics, commonsense, human fidelity, controllability
T2V-CompBench (sun2025t2vcompbench, ) Compositional T2V; tests binding of attributes/actions/relations with camera motion
Physical & spatial reasoning PhysBench (chow2025physbench, ) 10k video–image–text triplets; probes properties/relations/dynamics gaps in VLMs
UrbanVideo-Bench (zhao2025urbanvideo, ) 5.2k urban egocentric clips (16 tasks); recall, navigation, causal reasoning
Physics-IQ (motamed2025generative, ) Five domains (solids/fluids/optics/thermo/magnetism); law adherence vs. perceived realism
T2VPhysBench (guo2025t2vphysbench, ) Text-to-video; first-principles checklist of 12 core laws
VideoPhy (bansal2025videophy, ) Action-centric prompts; semantic adherence & physical commonsense; rule attribution
Basic Spatial Abilities (xu2025bsa, ) Psychometric framing; five spatial skills (perception/relations/orientation/rotation/visualization)
Embodied decision making EAI (li2024embodied, ) LLM-based agents; module-level (goal/subgoal/action/transition) evaluation; error taxonomy
EWMBench (yue2025ewmbench, ) AgiBotWorld manipulation/navigation; scene consistency, motion correctness, semantic alignment
WPE (quevedo2025evaluating, ) Policy evaluation in world model vs. real/sim; sequence-/part-level correspondence under identical actions
RoboScape (shang2025roboscape, ) Physics-informed embodied world model; policy lift & sim-to-real gap (data-engine role)
Benchmarking world models is both necessary and challenging. Because the community pursues divergent goals—learning internal representations vs. predicting future worlds—with heterogeneous technical approaches (e.g., LLM agents, video diffusion) and wide-ranging application domains (autonomous driving, robotics, social simulation), there is no single canonical task or metric. Nevertheless, several recent efforts illustrate how carefully designed testbeds can expose the specific gaps that prevent current models from becoming reliable world simulators, as summarized in Table 8.
Video-centric world simulation. WorldSimBench links perceptual quality to control by pairing human preferences with action-level consistency across sandbox, driving, and manipulation settings (qin2024worldsimbench, ); WorldScore complements this with a camera-specified protocol that decomposes performance into controllability, visual quality, and dynamics over 3,000 scenes, enabling head-to-head comparisons among 3D/4D and video generators (duan2025worldscore, ). VBench operationalizes largely automatic axes—temporal consistency, subject/background stability, and prompt adherence (huang2024vbench, )—while VBench-2.0 elevates intrinsic faithfulness (physics, commonsense, human fidelity, controllability) to separate “looking real” from “acting like a world” (zheng2025vbench, ). Compositional stress tests arrive via T2V-CompBench, which measures binding among attributes, actions, relations, and camera motion using MLLM-, detection-, and tracking-based metrics (sun2025t2vcompbench, ).
Physical and spatial reasoning. Beyond appearance, world models must respect physical law and support spatial competence. PhysBench (10k video–image–text triplets) reveals systematic gaps in object properties, relations, and dynamics for modern VLMs (chow2025physbench, ); UrbanVideo-Bench (5.2k drone clips, 16 task types) diagnoses deficits in recall, navigation, and causal reasoning in long egocentric streams (zhao2025urbanvideo, ). Law-centric suites sharpen the picture: Physics-IQ evaluates five domains (solid/fluid mechanics, optics, thermodynamics, magnetism), finding physical understanding largely decoupled from perceived realism (motamed2025generative, ); T2VPhysBench provides a first-principles checklist of 12 core laws for text-to-video systems (guo2025t2vphysbench, ). On the generative side, VideoPhy quantifies semantic adherence and physical commonsense under action-centric prompts, attributing errors to concrete rules (e.g., support, inertia, continuity) (bansal2025videophy, ). A complementary psychometric framing anchors five Basic Spatial Abilities—perception, relations, orientation, mental rotation, visualization—revealing geometry/rotation weaknesses across 13 VLMs and supplying calibrated tasks for progress tracking (xu2025bsa, ).
Embodied decision making. When world models are embedded in control loops, aggregate success rates obscure process failures. The Embodied Agent Interface (EAI) standardizes four LLM-based modules—goal interpretation, subgoal decomposition, action sequencing, transition modeling—and reports fine-grained error taxonomies (e.g., hallucination, affordance, planning) (li2024embodied, ). EWMBench tests embodied world models on scene consistency, motion correctness (physics-/task-consistent trajectories), and semantic alignment using AgiBot World data, explicitly tying video plausibility to action preconditions and affordances (yue2025ewmbench, ). A role-centered view evaluates models as environments: WPE compares rollouts under identical action sequences in the model versus real videos/simulators, reporting sequence- and part-level correspondences (quevedo2025evaluating, ); physics-informed world models such as RoboScape assess the data-engine role by measuring policy lift and sim-to-real gap when training on model-generated experience (shang2025roboscape, ).
@@ 7. Conclusion
Understanding the world and predicting the future have been long-standing objectives for scientists developing artificial generative intelligence, underscoring the significance of constructing world models across various domains. This paper presents the first comprehensive survey of world models that systematically explores their two primary functionalities: implicit representations and future predictions of the external world. We provide an extensive summary of existing research on these core functions, with particular emphasis on world models in decision-making, world knowledge learned by models, world models as video generation, and world models as embodied environments. Additionally, we review progress in key applications of world models, including generative games, robotics, autonomous driving, and social simulacra. Finally, recognizing the unresolved challenges in this rapidly evolving field, we highlight open problems and propose promising research directions with the hope of stimulating further investigation in this burgeoning area.
@@ Appendix B Figures and tables
Appendix C Update History
    2025.09.09: published version for ACM Computing Survey.
    2025.11.10: rewrite Sec. 2 (History and Current Development), summarize the roadmap of world models in the deep learning era (Fig. 2), reorganize application domains in Sec. 5, update recent papers.